For the following reaction: H2S + 2 KOH--- K2S + 2 H2O Calculate the normality of a solution that contains 3.80 grams of H2S in 784 mL of solution. ___N
Answer #1

Date molarity of solution = mole ay solute volume of salt (2) molar mass of H₂S = 2+32 = 349/mol moke cy HS 3.80 g = 0.111 m